15) yx+5)4 + 1 2 Ly 10 51 1051 10 x 5
Answer #1

y=-\frac{1}{2}\left(x+5\right)^4+1

to draw a graph points are

x y
-7 =-\frac{1}{2}\left(-7+5\right)^4+1=-\frac{1}{2}\left(-2\right)^4+1=-\frac{1}{2}(16)+1=-8+1=-7
-6 =-\frac{1}{2}\left(-6+5\right)^4+1=-\frac{1}{2}\left(-1\right)^4+1=-\frac{1}{2}(1)+1=-\frac{1}{2}+1=0.5
-5 =-\frac{1}{2}\left(-5+5\right)^4+1=-\frac{1}{2}\left(0\right)^4+1=-\frac{1}{2}(0)+1=0+1=1
-4 =-\frac{1}{2}\left(-4+5\right)^4+1=-\frac{1}{2}\left(1\right)^4+1=-\frac{1}{2}(1)+1=\frac{1}{2}+1=0.5
-3 =-\frac{1}{2}\left(-3+5\right)^4+1=-\frac{1}{2}\left(2\right)^4+1=-\frac{1}{2}(16)+1=-8+1=-7